User name R. Michaels
Log entry time 05:32:07 on October26,2001
Entry number 75039
This entry is a followup to: 75033
keyword=more tgt pump speed calib. Then no beam.
We keep trying to optimize and calibrate target pump speed.
The beam is back but unstable.
Run 1223.......TACH 750..... Oops, DT 25%. Stop to adjust ps factors
Run 1224.......TACH 750..... Not much data. Stopped due to RF problems.
I notice R-arm trigger 113 kHz, so expect some EDT (of order 1%).
Whew ! No beam.